Klapperich Lab:Notebook/Lab Meeting Notes/2008/12/16

From OpenWetWare

Jump to: navigation, search
Project name Main project page
Previous entry      Next entry

16 December 2008 Lab Meeting Agenda

  • Check with B&G about compressed air and vacuum?

† Announcements

  • No meeting on 12/23 or 12/30
  • Meeting Time for Spring Semester is THURSDAYS 3-5pm ROOM 705. Stating 1/8/09. See Lab Calendar for dates.


† Flu R01

  • First samples?
  • Plaque Assay.
  • Alex needs to be trained to make SU molds. Trained to make the monoliths (by Hussam By 1/15/09).


† SEPSIS Project

  • Sonali working on first draft of new version paper by 12/25. Journal?
  • Hussam working on Lambda phage data.


†RNA project

  • Report to Jeff from Cathie.
  • Do cDNA prep. Spec, gels, send to Jeff B. Shipping?

† COBRA

  • Mark- Control circuit - This will be done by January.
  • Evap quantification experiments ongoing. Still working on live/dead assay.(JD, JZ)

- problems: dead but green in the end
- called tech support.
- tests: supernatant fluorescent? excessive dye to observe time-lapse change in green/red ratio.
- omit washing step before dilution: wash, dye, incubate, dilute, instead of wash, dye, incubate, wash, dilute
- try running the experiment first, then running the live/dead dye through second.

  • Substrates (JZ)

- new set made, in glove box: the 1st 3 on the DOE. SERS signal strong, SEM today

  • DOE for the experiment.

- need to add humidity as it changes even in the glove box

  • Colloid experiment. inconclusive. Look at SEMs

- difficulty in locating the bacteria
- needs much higher particle/bacteria ratio
- experiments outside the channel first?

  • Prospectus framework/discussion: target the end of this week 1/23/09.


† Valve Array/Valve building (FJ)

  • Turn over platform to Teddy
  • TEDDY needs to meet with Mark.


† Fraunhofer: LOAC

  • So so close!


† Biointerfaces group


  • Strobe experiments? Jason cannot do them, perhaps MCK can. MCK will talk to Jason this week.
  • Made new design for trapper.
  • Mask writer was broken.


† CIMIT- Colson Grant

  • IRB still going/ revisions by tomorrow to Ian. Target date 1/1/09
  • STAR-CD Renewal - January ($2K)


† PCR

  • Improve the gel concentration 1.2% agarose. See if you can separate out primer dimers. UPDATE?
  • Look into Melting temp analysis on the 7300 - Done. what is contaminated?
  • Only 1/5 chips are good at the hot embossing step. The vaccum chuck will be checked and o-rings changed - helped a bit. ON HOLD

* Fixuture for PCR chip. The heater and the thermocouples would be on the fixture. ON HOLD

  • Try microscope for looking a fluor. Learn how to look at the data in Image J
  • 52, 60, 65 C repeats.


† RCA

  • Cathie submitted One pager Whitepaper. Waiting for response.


† Senior project

  • Megan 1)did some preliminary bonding with Zeonex and teflon film. 2) emailed off the Mask Design to FineLine

† F31: Cochlea

  • Set up adviser meeting Jan
  • Post Gent vs. Pre gent - 4 proteins. 3/4 no change.


† Silica Optimization (Lambda):

  • Update by 12/24.
  • Do three elutions.
  • Next E. coli.


Personal tools